OpenStudy (anonymous):

write as a single trig. fcn. cos2xcos3x - sin2xsin3x

OpenStudy (anonymous):

cos(a+b)=cos(a)cos(b) -sin(a)sin(b) so plug the values=cos(2x +3x)=cos(5x)